# Normal generators for mapping class groups

### v1

@@ 3. Lanier-Margalit’s well-suited criterion and its generalizations
The graph $N_{f}(S)$ is called graph of curves for $f$. The following is a slightly generalized version of Lanier-Margalit’s well-suited criterion:
@@
Let $f\in\operatorname{Mod}(S)$. If $N_{f}(S)$ is connected, then $\langle\langle f\rangle\rangle$ contains the commutator subgroup $[\operatorname{PMod}(S),\operatorname{PMod}(S)]$. Moreover, when $g\geq 3$, $N_{f}(S)$ is connected if and only if
@@
Since this is a neat criterion with many useful corollaries, we would like to include its proof here, which is organized slightly differently from [LM22].
@@
By Dehn-Lickorish’s theorem (Theorem 2.4), there exist finitely may non-separating simple closed curves $c_{1},\cdots,c_{k}$ on $S$ such that
@@
    for any $i,j\in\{1,\cdots,k\}$, the geometric intersection number $i(c_{i},c_{j})$ is either 00 or $1$;
@@
Step 1. We claim that the commutator subgroup $[\operatorname{PMod}(S),\operatorname{PMod}(S)]$ is contained in the normal closure of
@@
and hence
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]=\bigcup_{m\in\mathbb{N}}[P_{m},P_{m}].$
On the other hand, for $s,h,l\in\operatorname{PMod}(S)$, we observe
$[sh,l]=shlh^{-1}s^{-1}l^{-1}=s(hlh^{-1}l^{-1})s^{-1}(sls^{-1}l^{-1})=(s[h,l]s^{-1})[s,l]$
@@
$[P_{m},P_{m}]\leq\langle\langle[P_{m-1},P_{m-1}]\rangle\rangle.$
By an inductive argument, we conclude
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle[P_{1},P_{1}]\rangle\rangle.$
Moreover, for any $T_{c_{i}}$ and $T_{c_{j}}$, the commutator $[T_{c_{i}}^{\pm 1},T_{c_{j}}^{\pm 1}]$ is contained in the kernel of $\operatorname{PMod}(S)\to\operatorname{Mod}(S)/\langle\langle T_{c_{i}}T_{c_{j}}^{-1}\rangle\rangle$. This shows the claim.
@@
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle T_{c}T_{d}^{-1}\rangle\rangle.$
Fix two curves $c_{i},c_{j}$. Recall that $i(c_{i},c_{j})$ is either 00 or $1$. If $i(c_{i},c_{j})=0$, then $[T_{c_{i}},T_{c_{j}}]=e$. Otherwise, if $i(c_{i},c_{j})=1$, then it follows from $i(c,d)=1$ that there exists $h\in\operatorname{PMod}(S)$ such that $h(c)=c_{i}$ and $h(d)=c_{j}$. We then have
@@
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f\rangle\rangle.$
Indeed, we have $[\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle T_{c}T_{f(c)}^{-1}\rangle\rangle$ by Step 2. Since $T_{c}T_{f(c)}^{-1}=(T_{c}fT_{c}^{-1})f^{-1}\in\langle\langle f\rangle\rangle$, the claim follows.
@@
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f_{1}\cdots f_{m}\rangle\rangle.$
@@
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f\rangle\rangle$
@@
Step 5. For the last statements of the theorem, let us first assume that $g\geq 3$. By Harer’s theorem (Theorem 2.6), $\operatorname{PMod}(S)$ is perfect if $g\geq 3$, i.e., $[\operatorname{PMod}(S),\operatorname{PMod}(S)]=\operatorname{PMod}(S)$. Hence, in this case, if $N_{f}(S)$ is connected, then $\operatorname{PMod}(S)\leq\langle\langle f\rangle\rangle$ by Step 4. Since $\operatorname{PMod}(S)$ acts transitively on the set of non-separating curves, the converse easily follows. ∎
@@ 4. Small translation lengths and normal generation
Interestingly, whether a given mapping class is a normal generator can be detected by its dynamical behavior on Teichmüller space, as first shown by Lanier and Margalit [LM22] and extended by our joint work with Wu [BKW21b]. This section is devoted to the discussion on how dynamics on Teichmüller spaces is related to normal generation of mapping class groups.
@@
Let $S=S_{g,n}$ be a surface of genus $g\geq 1$ with $n\geq 0$ number of punctures. The Teichmüller space $\mathcal{T}(S)$ is defined as the set of marked hyperbolic structures on $S$. More precisely,
@@
There exists a natural metric $d_{\mathcal{T}}$ on $\mathcal{T}(S)$, called Teichmüller metric, which is invariant under the $\operatorname{Mod}(S)$-action. We omit the precise definition of the Teichmüller metric since we will not use it, while we focus more on some dynamical properties of the isometric $\operatorname{Mod}(S)$-action.
@@
Lanier and Margalit showed that pseudo-Anosovs with small asymptotic translation length on Teichmüller spaces are normal generators. More precisely, they showed the following:
Theorem 4.2.
@@
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f\rangle\rangle.$
@@
In fact, the original statement of Lanier and Margalit was about closed surface. However, the same proof works for surfaces with finitely many punctures, as the well-suited criterion (Theorem 3.1), which is a key ingredient, allows punctures. See [LM22, Section 3] for the related remark.
The pseudo-Anosov hypothesis on $f$ is crucial in the proof of Lanier-Margalit, as they investigated the combinatorics of a systole with respect to the singular Euclidean metric on the surface induced by a pseudo-Anosov, using [FLM08, Lemma 2.5, Proposition 2.7] which also work for punctured cases. Since the asymptotic translation length of a pseudo-Anosov mapping class is logarithmic of its stretch factor, as shown in Bers’ proof of Thurston’s classification theorem [Ber78], the translation length is captured by the singular Euclidean structure.
@@
Theorem 4.2 indeed implies the abundance of pseudo-Anosov normal generators, since there are pseudo-Anosov mapping classes whose asymptotic translation lengths arbitrary close to 00. More precisely, for each $g\geq 2$, let
[…]
where $S_{g}$ is a closed surface of genus $g$. Penner [Pen91] showed the asymptote of $L_{\mathcal{T}}(g)$ as follows:
@@
Therefore, for any closed surface of sufficiently large genus, we can always find a pseudo-Anosov mapping class whose asymptotic translation length on the Teichmüller space is smaller than $\frac{1}{2}\log 2$, which turns out to be a normal generator of the mapping class group by Theorem 4.2.
@@
Recall from Nielsen-Thurston classification (Theorem 2.5) that there are two types of infinite-order mapping classes, pseudo-Anosovs and reducibles. In our joint work with Wu [BKW21b], we extend Theorem 4.2 to deal with reducible mapping classes.
We are especially interested in reducible mapping classes with non-trivial dynamics. Let $f\in\operatorname{Mod}(S)$ be reducible. Then there are finitely many disjoint simple closed curves $c_{1},\cdots,c_{k}$ such that their union $\cup_{i=1}^{k}c_{i}$ is invariant under $f$. This implies that for some $m\in\mathbb{N}$, $f^{m}$ fixes each connected component of $S-\cup_{i=1}^{k}c_{i}$. Taking the collection $\{c_{1},\cdots,c_{k}\}$ to be maximal, restriction of $f^{m}$ to each component of $S-\cup_{i=1}^{k}c_{i}$ is either periodic or pseudo-Anosov. If every such restriction is periodic, then we can enlarge $m$ so that $f$ is a composition of Dehn twists $T_{c_{i}}$’s. It then follows that $\ell_{\mathcal{T}}(f)=0$, which makes considering the asymptotic translation length meaningless. In this regard, we come up with the following notion:
@@
Some terminologies similar to “partly pseudo-Anosov” were introduced by several authors. For instance, in [MM21], a mapping class is called “partial pseudo-Anosov” if its restriction to some subsurface is pseudo-Anosov while it is the identity outside of the subsurface. This is a more restrictive notion than partly pseudo-Anosov mapping classes. There is also a terminology “pure” mapping class, different from the pure mapping class that we use in this article, which has a similar feature to partly pseudo-Anosovs (see e.g. [BL23]).
We now state the generalizatin of Lanier-Margalit’s theorem (Theorem 4.2):
Theorem 4.6 (Baik-Kim-Wu).
@@
In fact, Theorem 4.8 still holds when $A$ has at least one genus and $S$ is of genus at least three. See [BKW21b, Theorem 3.1].
In the rest of this section, we describe the proof of Theorem 4.6, following [BKW21b]. From now on, suppose that $S$ is a closed surface. The key observation is that noraml generation of a mapping class can be detected by its local behavior:
@@
To prove Theorem 4.8, we first show the following lemma. Two disjoint non-separating simple closed curves $c,c^{\prime}$ are said to form a bounding pair if $S-(c\cup c^{\prime})$ is disconnected:
@@
Let $c,c^{\prime}$ are non-separating simple closed curves. Then there exists a sequence of simple closed curves $a_{0},\cdots,a_{k}$ such that
@@
  • if $a_{i-1}$ and $a_{i+1}$ are contained in the same connected componentn of $S-a_{i}$, then we can replace $a_{i}$ with a non-separating simple closed curve in the other component of $S-a_{i}$.
@@
We are now ready to prove the locality of normal generation. Let $f\in\operatorname{Mod}(S)$ and $A\subset S$ be as given. By well-suited criterion (Theorem 3.1), it suffices to show that $N_{f}(S)$ is connected.
Let $c,c^{\prime}$ be two vertices in $N_{f}(S)$, meaning that we take two non-separating simple closed curves. By Lemma 4.9, we may assume that $c,c^{\prime}$ are disjoint and do not form a bounding pair, to show that there exists a path in $N_{f}(S)$ between them.
Denoting by $g$ the genus of $S$, we have that $S-(c\cup c^{\prime})$ is of genus $g-2$ with four punctures, since $c$ and $c^{\prime}$ do not form a bounding pair. Since $A$ has genus at least three (in particular, at least two), there exists $h\in\operatorname{Mod}(S)$ such that $h(c)$ and $h(c^{\prime})$ are contained in $A$. Since $\operatorname{PMod}(A)$ acts transitively on the space of non-separating curves in $A$ and $\operatorname{PMod}(A)$ is contained in the normal closure of $f|_{A}$ in $\operatorname{Mod}(A)$, there exist $f_{1},\cdots,f_{n}\in\operatorname{Mod}(A)$ such that
@@
We then extend $f_{i}$’s to $S$ and obtain
@@
The last ingredient of the proof of Theorem 4.6 is about the structure of the Teichmüller space $\mathcal{T}(S)$. Let $c_{1},\cdots,c_{k}$ be a disjoint essential simple closed curves on $S$, which are not isotopic to each other. Then a marked hyperbolic structure on $S$ is determined by a restricted marked hyperbolic structure on $S-\cup_{i=1}^{k}c_{i}$, lengths of $c_{i}$’s, and how much it is twisted along $c_{i}$’s. For each $\sigma\in\mathcal{T}(S)$ and $c_{i}$, denote by $\ell_{\sigma}(c_{i})\in\mathbb{R}_{>0}$ the length of $c_{i}$ with respect to $\sigma$ and by $\tau_{\sigma}(c_{i})\in\mathbb{R}$ the angle that $\sigma$ is the angle twisted along $c_{i}$. Adding more curves if necessary, these quantities $\ell_{\sigma}$ and $\tau_{\sigma}$ parametrizes $\mathcal{T}(S)$, called Fenchel-Nielsen coordinates.
We set $\pi(\sigma)$ the restriction of $\sigma$ on $S-\cup_{i=1}^{k}c_{i}$. Using the upper half-plane model for the hyperbolic plane $\mathbb{H}^{2}=\{(x,y)\in\mathbb{R}^{2}:y>0\}$, we then have a map
@@
given by $\Pi(\sigma)=(\pi(\sigma),(\tau_{\sigma}(c_{1}),\ell_{\sigma}(c_{1})^{-1}),\cdots,(\tau_{\sigma}(c_{k}),\ell_{\sigma}(c_{k})^{-1}))$, where $\mathbb{H}_{i}^{2}$’s are copies of $\mathbb{H}^{2}$.
@@
Minsky studied how the Teichmüller metric $d_{\mathcal{T}}$ on $\mathcal{T}(S)$ and the metric $d_{X}$ on $X$ are related via the map $\Pi$ above [Min96]. He proved the following theorem, asserting that the region of $\mathcal{T}(S)$ on which the length of $\cup_{i=1}^{k}c_{i}$ is short has almost product structure in a metric sense. The surprising part is that there is only an additive error, not a multiplicative one (compare with a definition of a quasi-isometry):
@@
Let $f\in\operatorname{Mod}(S)$ and $A\subset S$ be as given. We consider $\partial A$ as the union of disjoint essential simple closed curves, which are not isotopic to each other. Fix $\varepsilon>0$ small enough so that Theorem 4.10 applies, and let $\delta>0$ be the constant given by it.
@@
and therefore $\operatorname{PMod}(A)$ is contained in the normal closure of $f|_{A}$ in $\operatorname{Mod}(A)$ by Theorem 4.2. By Theorem 4.8, we conclude

### v2

@@ 3. Lanier–Margalit’s well-suited criterion and its generalizations
The graph $N_{f}(S)$ is called graph of curves for $f$. Intuitively, if the normal closure of $f$ is large enough, then given a generic pair of non-separating essential simple closed curves, one can be mapped to the other by a sequence of conjugates of $f$. This means that the normal generation of $f\in\operatorname{Mod}(S)$ can be encoded by the connectivity of $N_{f}(S)$. Indeed, we have the following, which is a slight generalization of Lanier–Margalit’s well-suited criterion.
@@
Let $f\in\operatorname{Mod}(S)$. If $N_{f}(S)$ is connected, then $\langle\langle f\rangle\rangle$ contains the commutator subgroup $\operatorname{PMod}(S),\operatorname{PMod}(S)]$. Moreover, when $g\geq 3$, $N_{f}(S)$ is connected if and only if
@@
Since this is a neat criterion with many useful corollaries, we would like to include its proof here. The proof is organized slightly differently from [LM22].
@@
By Dehn–Lickorish’s theorem (Theorem 2.4), there exist finitely may non-separating simple closed curves $c_{1},\cdots,c_{k}$ on $S$ such that
@@
    for any $i,j\in\{1,\cdots,k\}$, the geometric intersection number $i(c_{i},c_{j})$ is either 0 or $1$;
@@
Step 1. We claim that the commutator subgroup $\operatorname{PMod}(S),\operatorname{PMod}(S)]$ is contained in the normal closure of
@@
hence
$\operatorname{PMod}(S),\operatorname{PMod}(S)]=\bigcup_{m\in\mathbb{N}}[P_{m},P_{m}].$
On the other hand, for $s,h,l\in\operatorname{PMod}(S)$, we observe that
$sh,l]=shlh^{-1}s^{-1}l^{-1}=s(hlh^{-1}l^{-1})s^{-1}(sls^{-1}l^{-1})=(s[h,l]s^{-1})[s,l]$
@@
$P_{m},P_{m}]\leq\langle\langle[P_{m-1},P_{m-1}]\rangle\rangle.$
By an inductive argument, we conclude that
$\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle[P_{1},P_{1}]\rangle\rangle.$
Moreover, for any $T_{c_{i}}$ and $T_{c_{j}}$, the commutator $T_{c_{i}}^{\pm 1},T_{c_{j}}^{\pm 1}]$ is contained in the kernel of $\operatorname{Mod}(S)\to\operatorname{Mod}(S)/\langle\langle T_{c_{i}}T_{c_{j}}^{-1}\rangle\rangle$, which is $\langle\langle T_{c_{i}}T_{c_{j}}\rangle\rangle$. Hence, the normal closure of the set $\{T_{c_{i}}T_{c_{j}}^{-1}:i,j\in\{1,\cdots,k\}\}$ contains all the commutators $T_{c_{i}}^{\pm 1},T_{c_{j}}^{\pm 1}]$, and therefore contains $P_{1},P_{1}]$. This shows the claim.
@@
$\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle T_{c}T_{d}^{-1}\rangle\rangle.$
Fix two curves $c_{i},c_{j}$. Recall that $i(c_{i},c_{j})$ is either 0 or $1$. If $i(c_{i},c_{j})=0$, then $T_{c_{i}},T_{c_{j}}]=e$. Otherwise, if $i(c_{i},c_{j})=1$, then it follows from $i(c,d)=1$ that there exists $h\in\operatorname{PMod}(S)$ such that $h(c)=c_{i}$ and $h(d)=c_{j}$. We then have
@@
$\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f\rangle\rangle.$
Indeed, we have $\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle T_{c}T_{f(c)}^{-1}\rangle\rangle$ by Step 2. Since $T_{c}T_{f(c)}^{-1}=(T_{c}fT_{c}^{-1})f^{-1}\in\langle\langle f\rangle\rangle$, the claim follows.
@@
$\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f_{1}\cdots f_{m}\rangle\rangle.$
@@
$\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f\rangle\rangle$
@@
Step 5. For the last statements of the theorem, let us first assume that $g\geq 3$. By Harer’s theorem (Theorem 2.6), $\operatorname{PMod}(S)$ is perfect if $g\geq 3$, i.e., $\operatorname{PMod}(S),\operatorname{PMod}(S)]=\operatorname{PMod}(S)$. Hence, in this case, if $N_{f}(S)$ is connected, then $\operatorname{PMod}(S)\leq\langle\langle f\rangle\rangle$ by Step 4. Since $\operatorname{PMod}(S)$ acts transitively on the set of non-separating curves, the converse easily follows. ∎
@@ 4. Small translation lengths and normal generation
Interestingly, whether a given mapping class is a normal generator can be detected by its dynamical behavior on Teichmüller space, as first shown by Lanier and Margalit [LM22] and extended by our joint work with Wu [BKW25]. This section is devoted to the discussion on how dynamics on Teichmüller spaces is related to normal generation of mapping class groups.
@@
Let $S=S_{g,n}$ be a surface of genus $g\geq 1$ with $n\geq 0$ punctures. The Teichmüller space $\mathcal{T}(S)$ is defined as the set of marked hyperbolic structures on $S$. More precisely,
@@
There exists a metric $d_{\mathcal{T}}$ on $\mathcal{T}(S)$, called Teichmüller metric, which is invariant under the $\operatorname{Mod}(S)$-action. We omit the precise definition of the Teichmüller metric since we will not use it, while we focus more on some dynamical properties of the isometric $\operatorname{Mod}(S)$-action.
@@
Lanier and Margalit showed that pseudo-Anosovs with small asymptotic translation lengths on Teichmüller spaces are normal generators. They exist due to Penner [PEN91] (see also Theorem 4.4). More precisely, they showed the following. We note that the upper bound $\frac{1}{2}\log 2$ does not depend on the surface $S$.
Theorem 4.2 (Lanier–Margalit’s criterion).
@@
$\operatorname{PMod}(S),\operatorname{PMod}(S)]\leq\langle\langle f\rangle\rangle.$
@@
In fact, the original statement of Lanier and Margalit was about closed surfaces. However, the same proof works for surfaces with finitely many punctures, as the well-suited criterion (Theorem 3.1), which is a key ingredient, allows punctures. See [LM22, Section 3] for the related remark.
The pseudo-Anosov hypothesis on $f$ is crucial in the proof of Lanier–Margalit, as they investigated the combinatorics of a systole with respect to the singular Euclidean metric on the surface induced by a pseudo-Anosov, using [FLM08, Lemma 2.5, Proposition 2.7] which also work for the punctured cases. Since the asymptotic translation length of a pseudo-Anosov mapping class is logarithmic in terms of its stretch factor, as shown in Bers’ proof of Thurston’s classification theorem [BER78], the translation length is captured by the singular Euclidean structure. Using this, Lanier–Margalit showed that $\ell_{\mathcal{T}}(f)\leq\frac{1}{2}\log 2$ implies the existence of an essential simple closed curve $c$ such that $i(c,f(c))\leq 2$, which was the key step in the proof of Theorem 4.2, providing a room to apply the well-suited criterion (Theorem 3.1).
@@
Theorem 4.2 indeed implies the abundance of pseudo-Anosov normal generators, since there are pseudo-Anosov mapping classes whose asymptotic translation lengths are arbitrary close to 0. More precisely, for each $g\geq 2$, let
(4.1) $L_{\mathcal{T}}(g):=\inf\{\ell_{\mathcal{T}}(f):f\in\operatorname{Mod}(S_{g})\text{ is pseudo-Anosov}\}$
where $S_{g}$ is a closed surface of genus $g$. Penner [PEN91] showed the following asymptotic behavior of $L_{\mathcal{T}}(g)$:
@@
Therefore, for any closed surface of sufficiently large genus, we can always find a pseudo-Anosov mapping class whose asymptotic translation length on the Teichmüller space is smaller than $\frac{1}{2}\log 2$. By Theorem 4.2, this mapping class turns out to be a normal generator of the mapping class group.
@@
Recall from Thurston’s classification (Theorem 2.5) that there are two types of infinite-order mapping classes, pseudo-Anosovs and reducibles. In our joint work with Wu [BKW25], we extend Theorem 4.2 to deal with reducible mapping classes.
We are especially interested in reducible mapping classes with non-trivial dynamics. Let $f\in\operatorname{Mod}(S)$ be reducible. Then there are finitely many disjoint simple closed curves $c_{1},\cdots,c_{k}$ whose union $\cup_{i=1}^{k}c_{i}$ is invariant under $f$. This implies that for some $m\in\mathbb{N}$, $f^{m}$ fixes each connected component of $S-\cup_{i=1}^{k}c_{i}$. Taking the collection $\{c_{1},\cdots,c_{k}\}$ to be maximal, the restriction of $f^{m}$ to each component of $S-\cup_{i=1}^{k}c_{i}$ is either periodic or pseudo-Anosov. If every such restriction is periodic, then we can enlarge $m$ so that $f$ is a composition of Dehn twists $T_{c_{i}}$’s. It then follows that $\ell_{\mathcal{T}}(f)=0$, which makes considering the asymptotic translation length meaningless. In this regard, we come up with the following notion:
@@
Some terminologies similar to “partly pseudo-Anosov” were introduced by several authors. For instance, in [MM21], a mapping class is called “partial pseudo-Anosov” if its restriction to some subsurface is pseudo-Anosov while it is the identity outside of the subsurface. This is a more restrictive notion than partly pseudo-Anosov mapping class. There is also a terminology “pure” mapping class, different from the pure mapping class that we use in this chapter, which has a similar feature to partly pseudo-Anosovs (see e.g. [BL23]).
We now state the generalizatin of Lanier–Margalit’s theorem (Theorem 4.2):
Theorem 4.6 (Baik–Kim–Wu).
@@
In fact, Theorem 4.8 still holds when AA has genus $\geq 1$ and $S$ is of genus at least three. See [BKW25, Theorem 3.1].
In the rest of this section, we describe the proof of Theorem 4.6, following [BKW25]. From now on, suppose that $S$ is a closed surface. The key observation is that normal generation of a mapping class can be detected by its local behavior:
@@
To prove Theorem 4.8, we first show the following lemma. Two disjoint non-separating simple closed curves $c,c^{\prime}$ are said to form a bounding pair if $S-(c\cup c^{\prime})$ is disconnected.
@@
Let $c,c^{\prime}$ be non-separating simple closed curves. Then there exists a sequence of simple closed curves $a_{0},\cdots,a_{k}$ such that
@@
  • if $a_{i-1}$ and $a_{i+1}$ are contained in the same connected component of $S-a_{i}$, then we can replace $a_{i}$ with a non-separating simple closed curve in the other component of $S-a_{i}$.
@@
We are now ready to prove the locality of normal generation. Let $f\in\operatorname{Mod}(S)$ and $A\subset S$ be as given. By a well-suited criterion (Theorem 3.1), it suffices to show that $N_{f}(S)$ is connected.
Let $c,c^{\prime}$ be two vertices in $N_{f}(S)$. In other words, we take two non-separating simple closed curves. By Lemma 4.9, we may assume that $c,c^{\prime}$ are disjoint and do not form a bounding pair, to show that there exists a path in $N_{f}(S)$ between them.
Denoting by $g$ the genus of $S$, we have that $S-(c\cup c^{\prime})$ is of genus g2g-2 with four punctures, since $c$ and $c^{\prime}$ do not form a bounding pair. Since $A$ has genus at least three (in particular, at least two), there exists $h\in\operatorname{Mod}(S)$ such that $h(c)$ and $h(c^{\prime})$ are contained in $A$. Since $\operatorname{PMod}(A)$ acts transitively on the space of non-separating curves in $A$ and $\operatorname{PMod}(A)$ is contained in the normal closure of $f|_{A}$ in $\operatorname{Mod}(A)$, there exist $f_{1},\cdots,f_{n}\in\operatorname{Mod}(A)$ such that
@@
We then extend the $f_{i}$’s to $S$ and obtain
@@
The last ingredient of the proof of Theorem 4.6 is about the structure of the Teichmüller space $\mathcal{T}(S)$. Let $c_{1},\cdots,c_{k}$ be a disjoint essential simple closed curves on $S$ which are not isotopic to each other. Then a marked hyperbolic structure on $S$ is determined by a restricted marked hyperbolic structure on $S-\cup_{i=1}^{k}c_{i}$, lengths of $c_{i}$’s, and how much it is twisted along $c_{i}$’s. and for each $c_{i}$, the length and the amount of twisting along $c_{i}$. For each $\sigma\in\mathcal{T}(S)$ and $c_{i}$, denote by $\ell_{\sigma}(c_{i})\in\mathbb{R}_{>0}$ the length of $c_{i}$ with respect to $\sigma$ and by $\tau_{\sigma}(c_{i})\in\mathbb{R}$ the angle that $\sigma$ is the angle twisted along $c_{i}$. Adding more curves if necessary, these quantities $\ell_{\sigma}$ and $\tau_{\sigma}$ parametrize $\mathcal{T}(S)$. This parametrization is called Fenchel–Nielsen parametrization.
We set $\pi(\sigma)$ to be the restriction of $\sigma$ on $S-\cup_{i=1}^{k}c_{i}$. Using the upper half-plane model for the hyperbolic plane $\mathbb{H}^{2}=\{(x,y)\in\mathbb{R}^{2}:y>0\}$, we then have a map
@@
given by $\Pi(\sigma)=(\pi(\sigma),(\tau_{\sigma}(c_{1}),\ell_{\sigma}(c_{1})^{-1}),\cdots,(\tau_{\sigma}(c_{k}),\ell_{\sigma}(c_{k})^{-1}))$, where the $\mathbb{H}_{i}^{2}$’s are copies of $\mathbb{H}^{2}$.
@@
Minsky studied how the Teichmüller metric $d_{\mathcal{T}}$ on $\mathcal{T}(S)$ and the metric $d_{X}$ on $X$ are related via the map $\Pi$ [MIN96]. He proved the following theorem, asserting that the region of $\mathcal{T}(S)$ on which the length of $\cup_{i=1}^{k}c_{i}$ is short has an almost product structure in a metric sense. The surprising part is that there is only an additive error, not a multiplicative one (compare with a definition of a quasi-isometry):
@@
Let $f\in\operatorname{Mod}(S)$ and $A\subset S$ be as given. We consider $\partial A$ as the union of disjoint essential simple closed curves which are not isotopic to each other. Fix $\varepsilon>0$ small enough so that Theorem 4.10 applies, and let $\delta>0$ be the constant given by it.
@@
therefore by Theorem 4.2, $\operatorname{PMod}(A)$ is contained in the normal closure of $f|_{A}$ in $\operatorname{Mod}(A)$. By Theorem 4.8, we conclude that